Part a) In a nitrous acid-nitrite buffer, the pH is 2 pH units lower than pKa of the acid. If the concentration of nitrite is 0.45M, calculate the concentration of the conjugate acid.
Part b) determine if the reaction of this conjugate acid/base is spontaneous or not. (Use delta G to figure this out)
1
Expert's answer
2017-03-20T09:43:05-0400
a) pH = pKa + lg(Cb/Ca) so Ca = Cb/(10^(pH - pKa)) = 4.5*10-3 M.

b) G = -RTlg(Ka) = -8.31J/mol*K*273K*(-7.7) = 17.7 kJ/mol (not spontaneous)